OpenStudy (exoexoexo):

if 15 cans of food are needed for 7 men for 2 days ], the number of cans needed for 4 men for 7 days is?

OpenStudy (anonymous):

total cans per day = 15/2 (since 2 days) then /7 for per adult per day = 15/14 cans per adult per day Then multiply by # adults * # days 15/14 * 4 * 7 = 15/14 * 28 = 15*2 = 30 cans required

OpenStudy (anonymous):

@EXOEXOEXO

OpenStudy (exoexoexo):

thank you

OpenStudy (anonymous):

you're welcome :)